# HEIGHT: HEterogeneous Interaction GrapH Transformer for Robot Navigation in Crowded and Constrained Environments

### v2

@@ V-C1 Effectiveness of scene representation
Similarly, by representing obstacles as groups of circles, DS-RNN performs better in the More crowded environment (Fig. 7(d)) than in the More constrained environment (Fig. 6(d)). In addition, Table II shows that DS-RNN achieves the highest average collision rate with obstacles (0.13) in all environments. Furthermore, the obstacle collision rate of DS-RNN increases in all 4 OOD environments, indicating an overfitting problem. Among the OOD environment, the percentage of obstacle collision increase is higher in environments with higher obstacle-to-human ratios. For example, in Less crowded with the highest percentage of obstacles, the obstacle collision
@@
rate (0.21) increases by 133% w.r.t. the obstacle collision rate in training distribution (0.09). On the contrary, in More crowded with the lowest percentage of obstacles, the obstacle collision rate (0.06) drops by 33%. The reason is that DS-RNN has trouble inferring the geometric shapes of obstacles from a large group of circles, and thus fails to avoid collision with them. Thus, treating both humans and obstacles as circles is not an ideal input representation for robot navigation algorithms.
@@
By splitting human and obstacle input representations, HomoGAT and HEIGHT achieve the top 2 performance across most metrics and environments. Especially, HEIGHT learns to keep a larger distance from human paths yet a shorter distance from obstacles to balance safety and efficiency (Fig. 6(g)), indicated by the lowest overall collision rates (0.17) and the shortest average navigation time (18.10) in all environments. Therefore, we conclude that our split representation consisting of detected human states and obstacle point clouds is most suitable to learn robot navigation in crowded and constrained environments with RL, among all popular choices in Table II.
V-C2 Effectiveness of the heterogeneous st-graph
To justify our heterogeneous st-graph, we compare our method with $\textrm{A}^{*}$+ CNN and DRL-VO with no graph concept and HomoGAT with a homogenous graph in Table II.
@@ V-C3 Importance of attention networks
We use ablations to evaluate the contribution of RH and HH attention networks to the performance of HEIGHT, as shown in Table III. A visualization of attention weights can be found in Fig. 10 in Appendix.
First, in terms of all metrics in most environments, the model with both RH and HH attention shows the strongest results, followed by the models with only one attention, and finally by the model with no attention. For example, in Fig. 7, the ablations in (a), (b), and (c) fail yet our full model in (e) succeeds.Especially in the 2 most challenging More crowded and More constrained environments, the existence of either RH or HH attention significantly boosts the average success rate, compared with No attn. For example, in Fig. 6 (e), the robot detours and goes dangerously close to the crowds in a constrained area, yet HH attention allows it to realize the danger and resume to a less crowded path. This result shows that reasoning about both HH and RH spatial relationships is essential for our problem, especially in more crowded or more constrained environments where the spatial interactions are also dense.
Second, by comparing RH and HH, we find that HH attention plays a more important role in all environments in terms of success rate and navigation time. This is because the number of HH edges is larger than the number of RH edges in most cases. Thus, the robot can observe and is affected by a relatively larger number of HH interactions yet a smaller number of RH interactions.
@@ V-C4 Failure cases
By visualizing the testing episodes across all environments, we find that HEIGHT typically collides when (1) a human arrives at its goal and suddenly changes directions to a new goal (Fig. 11(a) in Appendix) or (2) the robot surroundings are extremely crowded and constrained, and almost all free paths are blocked by humans (Fig. 11(b)). In these cases, due to its speed limit, the robot sometimes cannot switch to an alternative path in time to prevent collisions. To remedy the difficulty of RL in long-term decision making [61], in future work, our method can be combined with long-term prediction and path planning algorithms that consider the stochasticity of pedestrian motions.
Besides answering the four research questions in Sec. V, we provide additional insights obtained from simulation ex- periments in Appendix-C.
Figure 8: A testing episode of our method in the real Hallway environment. The blue arrow denotes the robot path that results from its actions. The red star denotes the goal position. In this episode, the turtlebot avoids two pedestrians, one after another in a narrow corridor, enters a narrow doorway, and arrives at the goal. More qualitative results can be found in the video attachment.
@@ VI Real-world Experiments
In this section, we present our hardware setup and sim2real testing results in everyday environments with pedestrians and static constraints.
VI-A Experiment setup
We train the sim2real policies in the simulators of a hallway (Fig. 5(b)) and a lounge (Fig. 5(c)) for $2\times 10^{8}$ timesteps with a decaying learning rate $5\times 10^{-5}$. To learn robust policies, we inject noises into the agent positions and robot control. Then, as shown in Fig. 5 (d) and (e), we test the policies in the two corresponding everyday environments in a university building without any additional training in the real-world. In the hallway environment where the free space is extremely narrow, we test the robot’s ability to handle constraints with low density crowds. In the lounge environment, we test the robot’s ability to avoid dense crowds and obstacles with more diverse shapes. Details about pedestrians and hardware configuration can be found at Appendix -D.
Our baseline is the ROS navigation stack, which uses the dynamic window approach (DWA) [1] as the local planner and $\text{A}^{*}$ as the global planner. For each method, we run 30 trials in total. Among all trials, the start goal positions of the robot are the same, while the number and trajectories of pedestrians are similar. We measure the success, collision, timeout rates, and navigation time of successful episodes as testing metrics.
TABLE IV: Real-world results in 2 everyday environments
@@ VI-A Experiment setup
Hallway 1-2 humans 12 Navigation Stack 0.72 0.06 0.22 16.71
@@
Lounge 1-6 humans 18 Navigation Stack 0.83 0.17 0.00 32.00
@@
VI-B Results
In the highly constrained yet less crowded Hallway environment, ROS navigation stack often needs to spin in place to replan, as shown by the higher navigation time in Table IV. Some of the spinning recovery attempts fail and result in timeouts. In ROS navigation stack, both global and local planners treat humans as obstacles. As a result, similar to the baselines in Sec. V-C1, the robot has difficulties distinguishing dynamic obstacles that will clear out and static obstacles that will always stay, causing failures in highly constrained spaces. On the contrary, by using different representations of humans and obstacles, our method is able to explore the different strategies to avoid them through trial and error during RL training. Consequently, the robot demonstrates high success rates in Table IV. Qualitatively, the robot is able to safely pass a human in a narrow corridor (Fig. 8 (a)-(d)), takes a wide turn to give a human enough room, and enters an extremely narrow doorway (Fig. 8 (e)-(h)).
In the less constrained yet more crowded Lounge environment, ROS navigation stack and ours have the same success and collision rate, because the Lounge environment has enough space for the navigation stack to apply a “stop and wait” strategy for humans and obstacles. However, as a cost, “stop and wait” takes a longer average time to arrive at goals as shown in the last column of Table IV. Different from this naïve strategy, HEIGHT adapts its navigation behaviors based on different types of spatio-temporal interactions. For example, in Fig. 9 (a)-(c), since pedestrians walking in opposite directions are crossing each other, the robot first turns left to avoid the lady on its right, and then turns right to avoid the two males on its left. Then, in Fig. 9 (d)-(e) and (f)-(g), the robot chooses actions that deviate from the current and intended paths of pedestrians, walls, and the table to safely arrive at the goal. The failure cases of both methods are caused by the tables with irregular 3D shapes. Due to the height of the 2D LiDAR, some wide parts of the tables are not detected or mapped, and thus the robot cannot avoid them. Further pre-processing of the LiDAR point clouds or adding 3D sensors can mitigate this problem.
Furthermore, the successful sim2real transfer from a cheap and low-fidelity simulator to complex everyday environments demonstrates the robustness of our input representation and the cost-efficiency of our pipeline design.
@@ VII Discussions, Limitations, and Future Work
Deep RL is a promising tool to solve robotic problems that are beyond the capabilities of traditional rule-based methods without large-scale real-world datasets. However, preventing the performance degradation of a deep RL policy when inevitable distribution shifts happen, especially in real-world, is challenging. In this section, we reflect on the key components of our framework, discuss the limitations of our approach, and propose directions for future research.
@@ VII-A Sim2real through Real2sim
Lessons learned: To overcome sim2real gaps, the design of simulation pipelines needs to be guided by the constraints of hardware and environments in the real-world. On one hand, we determine the input representation of HEIGHT based on what could be obtained from sensors and off-the-shelf perception modules and how accurate they are. We find that intermediate features, such as detected human positions and processed point clouds, reduces sim2real gaps. On the other hand, we also ensure the consistency of the simulation and real-world, such as the robot action space, whenever possible. These design choices allow our policy to generalize to different simulation environments and deploy to challenging real-world scenarios.
Limitation and future work: However, although we have minimized the sim2real gaps through real2sim, certain gaps still exist. Before sim2real transfer, the policy needs to be finetined in simulations that mimic the target environments (Fig.5(d) and (e)). Despite this, in real-world experiments, the difficulty of the task is reduced and the agility of the robot policy is only partially transferred from simulation. To further align the simulation and the real-world, we plan to explore the following directions for future work: (1) developing a more natural pedestrian model to replace the ORCA humans in the simulator, (2) revising our pipeline to enable failure detections and self-supervised RL fine-tuning in the real-world [62, 63], (3) using a small amount of real-world data to automatically optimize the parameters of our simulator to match the real-world environment [64, 65].
@@ VII-B Scene representation
Lessons learned: A good scene representation should be tailored to the needs of its downstream task. Besides the above sim2real considerations, our scene representation is split due the different nature of humans and obstacles for robot collision avoidance. The size of humans are small and their shapes are simple. Therefore, to avoid humans, the robot only needs to treat them as circles with a heading direction. In contrast, obstacles have larger and more complicated surfaces. The part of the obstacle contours that faces the robot is more important for collision avoidance. Therefore, point clouds are the most intuitive way to represent such useful information about obstacles. Our experiments in Sec. I show that this split scene representation reduces robot collision avoidance with both dynamic and static obstacles.
Limitation and future work: A side effect of our scene abstraction is the loss of detailed information such as gaits of humans and 3D shapes of obstacles. However, we argue that this is a tradeoff to minimize sim2real gaps with limited simulation tools and computation resources. Another side effect is the cascading errors between the perception modules and robot policy, such as inaccurate human detections or robot localization. A future direction to automate the design of inputs and outputs is the joint optimization of the whole robotic stack from perception to control [66, 67].
@@ VII-C Structured neural network
Limitation and future work: The structured neural network utilizes the inductive bias behind human behaviors to solve specific problems. However, without re-training, the aforementioned network may not generalize to other human-robot interaction (HRI) scenarios such as autonomous driving. In other fields such as computer vision and natural language processing, foundation models that are trained on large-scale diverse data have shown promise in generalization. Inspired by the standardized benchmarks of autonomous driving and robot manipulation [68, 69], we aim to contribute to the benchmarking efforts of human-centered tasks.
@@ VII-D Training method
Lessons learned: Deep RL enables the robot to explore the environment and learn meaningful behaviors through trial and error. Without heuristics or demonstrations, the robot has to collect reward signals to improve its policy. Consequently, simulation development with a randomization scheme and a good reward design are indispensable to the performance of our method.
Limitation and future work: Nevertheless, our method is subjective to the inherent limitations of RL, such as the need to build simulations, low training data efficiency, and difficulties in long-horizon tasks. In future work, combining the RL policy with other traditional planners or imitation learning has the potential to alleviate these problems [70].

### v3

@@ V-C1 Effectiveness of scene representation
Similarly, by representing obstacles as groups of circles, DS-RNN performs better in the More crowded environment (Fig. 6(d)) than in the More constrained environment (Fig. 5(d)). In addition, Table II shows that DS-RNN achieves the highest average collision rate with obstacles (0.13) in all environments. Furthermore, the obstacle collision rate of DS-RNN increases in all 4 OOD environments, indicating an overfitting problem. Among the OOD environment, the percentage of obstacle collision increase is higher in environments with higher obstacle-to-human ratios. For example, in Less crowded with the highest percentage of obstacles, the obstacle collision rate (0.21) increases by 133% w.r.t. the obstacle collision rate in training distribution (0.09). On the contrary, in More crowded with the lowest percentage of obstacles, the obstacle collision rate (0.06) drops by 33%. The reason is that DS-RNN has trouble inferring the geometric shapes of obstacles from a large group of circles, and thus fails to avoid collision with them. Thus, treating both humans and obstacles as circles is not an ideal input representation for robot navigation algorithms.
By splitting human and obstacle input representations, HomoGAT and HEIGHT achieve the top 2 performance across most metrics and environments. Especially, HEIGHT learns to keep a larger distance from human paths yet a shorter distance from obstacles to balance safety and efficiency (Fig. 5(g)), indicated by the lowest overall collision rates (0.17) and the shortest average navigation time (18.10) in all environments. In addition, in the last 3 frames of Fig. 6(e), the robot detours around the thin and long wall because the wall will not move out of its way. However, when the robot encounters the leftmost green human, it waits for the human to pass and resume afterwards because the human only blocks its way temporarily. Therefore, we conclude that our split representation consisting of detected human states and obstacle point clouds is most suitable to learn robot navigation in crowded and constrained environments with RL, among all popular choices in Table II.
V-C2 Effectiveness of the heterogeneous st-graph
To justify our heterogeneous st-graph, we compare our method with ORCA, $\textrm{A}^{*}$+ CNN, and DRL-VO with no graph concept and HomoGAT with a homogenous graph in Table II.
ORCA models the RH and OA interactions at the current timestep as velocity obstacles. As a result, ORCA is the second worst among all methods in terms of success rate and performs especially poorly in terms of obstacle collision rates, shown in Fig. 5(g). Since all methods that outperform ORCA contain either a search-based planner or recurrent networks, this result indicates the importance of HH interaction and long-term temporal reasoning in our problem. However, ORCA has the shortest average path length in most environments due to its greedy decisions to move toward the goal.
@@ V-C3 Importance of attention networks
We use ablations to evaluate the contribution of RH and HH attention networks to the performance of HEIGHT, as shown in Table III. A visualization of attention weights is in Fig. 10 in Appendix.
First, in terms of all metrics in most environments, the model with both RH and HH attention shows the strongest results, followed by the models with only one attention, and finally by the model with no attention. For example, in Fig. 6, the ablations in (a), (b), and (c) fail yet our full model in (e) succeeds. Especially in the 2 most challenging More crowded and More constrained environments, the existence of either RH or HH attention significantly boosts the average success rate by 0.38 and 0.43 respectively, compared with No attn. For example, in Fig. 5 (e), the robot detours and goes dangerously close to the crowds in a constrained area, yet HH attention allows it to realize the danger and resume to a less crowded path. This result shows that reasoning about both HH and RH spatial relationships is essential for our problem, especially in more crowded or more constrained environments where the spatial interactions are also dense.
Second, by comparing RH and HH, we find that HH attention plays a more important role in all environments in terms of success rate and navigation time. An extreme example can be observed from more constrained environment, where the success rate of RH+HH and HH are the same.
This is because the number of HH edges is larger than the number of RH edges in most cases. Thus, the robot can observe and is affected by a relatively larger number of HH interactions yet a smaller number of RH interactions.
@@ V-C4 Failure cases
By visualizing the testing episodes across all environments, we find that HEIGHT typically collides when (1) a human arrives at its goal and suddenly changes directions to a new goal (Fig. 11(a) in Appendix) or (2) the robot surroundings are extremely crowded and constrained, and almost all free paths are blocked by humans (Fig. 11(b)). In these cases, due to its speed limit, the robot sometimes cannot switch to an alternative path in time to prevent collisions. To remedy the difficulty of RL in long-term decision making, in future work, our method can be combined with motion primitives and path planning algorithms that consider the stochasticity of pedestrian motions [hu2025composablenav].
Besides answering the four research questions in Sec. V, we provide additional insights in Appendix-C.
[…]
Figure 8: A deployment scenario of our method in an outdoor environment with 35+ humans. The jackal takes detours if there is enough space, decelerates if all paths are blocked, and merges into the traffic that moves in the same direction if possible.
@@ VI Real-world Experiments
In this section, we present challenges, setup, and results in everyday environments in the real-world.
VI-A Setup
Challenges and solutions: Real-world navigation with pedestrians and static obstacles is challenging because perception errors, unexpected human behaviors, and noisy robot dynamics model cause sim2real gaps. To deal with noises in human detection, robot localization, and robot dynamics, we inject Gaussian noises into the observed agent positions and the robot dynamics model in the simulator to robustify the learned policy. To conquer the human behavior gap, in the simulator, we design pedestrian routes based on observations of human flow and the layout of the deployment environment.
We test the policies in four everyday environments without any additional training in the real-world. The hallway environment (Fig. 12) has extremely narrow free space and tests the robot’s ability to handle static obstacles with low density crowds. The lounge environment (Fig. 13) tests the robot’s ability to avoid more pedestrians and obstacles with more diverse shapes. The atrium (Fig. 7) and the outdoor environment (Fig. 8) test the robot’s potential of real-world deployment in dense and unrehearsed crowds. Details about setup can be found in Appendix-D.
In the Hallway and Lounge environments, we conduct controlled experiments where the trajectories of all pedestrians are fixed to ensure a fair comparison. We compare our method with the following two baselines: (1) ROS navigation stack is a geometric planner which uses the dynamic window approach (DWA) [fox1997dynamic] as the local planner and $\text{A}^{*}$ as the global planner; (2) DRL-VO is a state-of-the-art RL method [xie2023drlvo]. For each method, we run 30 trials in total. Among all trials, the start goal positions of the robot are fixed. We measure the success, collision, timeout rates, and navigation time of successful episodes as testing metrics.
To demonstrate the real-world deployment of our method, we conduct in-the-wild experiments in two densely packed public spaces: the Atrium and Outdoor environment in Fig. 7 and Fig. 8 with non-rehearsed pedestrian flows.
VI-B Baseline comparison results
In the highly constrained yet less crowded Hallway environment, ROS navigation stack often needs to spin in place to replan, as shown by the higher navigation time in Table IV. Some of the spinning recovery attempts fail and result in timeouts. In ROS navigation stack, both global and local planners treat humans as obstacles. As a result, similar to the baselines in Sec. V-C1, the robot has difficulties distinguishing dynamic obstacles that will clear out and static obstacles that will always stay, causing failures in highly constrained spaces. For the same reason as simulation experiment in Sec. V-C-1), the large input representation and network of DRL-VO pose challenges for RL exploration in this highly constrained environment. Insufficient exploration leads to an underfitting problem, where the robot collides with obstacles or goes back and forth without arriving at the goal. On the contrary, by using different representations of humans and obstacles, our method is able to explore the different strategies to avoid them through trial and error during RL training. For example, in Fig. 12(e)-(g), when turning into the narrow doorway, the robot takes a wide turn to yield enough space so that the human can exit the doorway. However, the robot did not exhibit such prediction-aware behaviors when avoiding obstacles. Consequently, the robot demonstrates high success rates in Table IV. Qualitatively, the robot is able to safely pass a human in a narrow corridor (Fig. 12 (a)-(d) in Appendix), takes a wide turn to give a human enough room, and enters an extremely narrow doorway (Fig. 12 (e)-(h)).
TABLE IV: Real-world baseline comparison results
@@ VI-B Baseline comparison results
Hallway 1-2 humans 18 Navigation Stack 0.72 0.06 0.22 16.71
DRL-VO 0.50 0.50 0.00 29.33
@@
Lounge 1-6 humans 12 Navigation Stack 0.83 0.17 0.00 32.00
DRL-VO 0.67 0.33 0.00 28.91
@@
In the less constrained yet more crowded Lounge environment, DRL-VO has the highest collision rate due to its underfitting problem. In addition, since only DRL-VO requires human velocities as inputs, the noises of human velocity estimation exaggerates its problem. ROS navigation stack and ours have the same success and collision rate, because the Lounge environment has enough space for the navigation stack to apply a “stop and wait” strategy for humans and obstacles. However, as a cost, “stop and wait” takes a longer average time to arrive at goals as shown in the last column of Table IV. Different from this naïve strategy, HEIGHT adapts its navigation behaviors based on different types of spatio-temporal interactions. For example, in Fig. 13 (a)-(c) in Appendix, since pedestrians walking in opposite directions are crossing each other, the robot first turns left to avoid the lady on its right, and then turns right to avoid the two males on its left. Then, in Fig. 13 (d)-(e) and (f)-(g), the robot chooses actions that deviate from the current and intended paths of pedestrians, walls, and the table to safely arrive at the goal. The failure cases of both methods are caused by changes in the arrangement of tables and chairs, while the fake point clouds are generated from a map with outdated furniture positions. This issue can be mitigated by replacing fake point clouds with real-time point clouds from sensors.
VI-C Deployment results
The behavior mismatch between ORCA humans in simulation and real pedestrians is exaggerated during deployment. Since real pedestrians are usually risk-averse and conservative when interacting with the robot, all simulated humans react to the robot with a larger safety space. If the simulated humans out of robot’s view are about to collide with the robot, they freeze until they are far enough from the robot. To conquer perception and dynamics gaps, we strictly duplicate ZED 2i camera’s sensor range (120 field-of-view (FoV) and humans are only detectable $\sim1$5 meters from the camera) and use a kinematic bicycle model with measured delays in simulation.
As shown in Fig. 7 and Fig. 8, during deployment, our method has learned an adaptive policy that accommodates the robot’s limitations. When the crowd is sparse, the robot circles around pedestrians with a large safety distance to keep them in its FoV as long as possible. When the crowd is dense, the robot stops temporarily and keeps adjusting its orientation to find a direction with enough space to move forward, and resumes right after the pedestrians pass the robot. In the outdoor environment where pedestrians form a two-way traffic, the robot shows a smart strategy by merging into the traffic that moves in the same direction. These observations demonstrate that our method has learned an adaptive policy that is robust to sim2real gaps and can navigate safely and efficiently with unpredictable humans.
@@ VII Discussions, Limitations, and Future Work
In this section, we reflect on the key components of our framework, discuss the limitations of our approach, and propose directions for future research.
@@ VII-A Sim2real through Real2sim
Lessons learned: To overcome sim2real gaps, the design of simulation pipelines needs to be guided by the constraints of hardware and environments in the real-world. On one hand, we determine the input representation of HEIGHT based on what could be obtained from sensors and perception modules and how accurate they are. We find that intermediate features, such as detected human positions and processed point clouds, reduce sim2real gaps. On the other hand, we also ensure the consistency of the simulation and real-world, such as the robot dynamics model, whenever possible. These design choices allow our policy to generalize to different simulation environments and be deployed in challenging real-world scenarios.
Limitation and future work: Although we have minimized the sim2real gaps through real2sim, certain gaps still exist. Before sim2real transfer, the policy needs to be finetined in simulations that mimic the target environments (Fig.9). To further align the simulation and the real-world, we plan to explore the following directions for future work: (1) developing a more natural pedestrian model to replace the ORCA humans in the simulator, (2) revising our pipeline to enable failure detections and self-supervised RL fine-tuning in the real-world [Zhu2020The, chang2023data], (3) using a small amount of real-world data to automatically optimize the parameters of our simulator to match the real-world environment [du2021auto, lim2022real2sim2real].
@@ VII-B Scene representation
Lessons learned: A good scene representation should be tailored to the needs of its downstream task. Besides the above sim2real considerations, our scene representation is split due the different nature of humans and obstacles for robot collision avoidance. The size of humans are small and their shapes are simple. Therefore, to avoid humans, the robot only needs to treat them as moving circles. In contrast, obstacles have larger and more complicated shapes. The part of the obstacles that faces the robot is more important for collision avoidance. Therefore, point clouds are the most intuitive way to represent obstacles. Our experiments in Sec. V show that this split scene representation reduces robot collision avoidance with both dynamic and static obstacles.
Limitation and future work: A side effect of our scene abstraction is the loss of detailed information such as gaits of humans and 3D shapes of obstacles. Another side effect is the cascading errors between the perception modules and robot policy, such as inaccurate human detections or robot localization. A future direction is the joint optimization of the whole robotic stack from perception to control [hu2023_uniad, wang2024mosaic].
@@ VII-C Structured neural network
Limitation and future work: The structured neural network utilizes the inductive bias behind human behaviors to solve specific problems. However, without re-training, the network cannot generalize to other tasks such as autonomous driving. In other fields such as computer vision and natural language processing, foundation models that are trained on large-scale diverse data have shown promise in generalization. Inspired by the standardized benchmarks of autonomous driving and robot manipulation [caesar2021nuplan, openxembodiment2023], we aim to contribute to the benchmarking efforts of human-centered tasks.
@@ VII-D Training method
Lessons learned: Deep RL is a promising tool to solve robotic problems that are beyond the capabilities of traditional rule-based methods without large-scale real-world datasets. It enables the robot to explore the environment and learn meaningful behaviors through trial and error. Without heuristics or demonstrations, the robot has to collect reward signals to improve its policy. Consequently, simulation development with a randomization scheme and a good reward design are indispensable to the performance of our method.
Limitation and future work: Our method is subjective to the inherent limitations of RL, such as the need for simulators, low training data efficiency, and difficulties in long-horizon tasks. In addition, preventing performance degradation of the policy when inevitable distribution shifts happen, especially in real-world, is challenging. In future work, combining the RL policy with other planners, imitation learning, or foundation models may alleviate these problems [raj2024rethinking].
VII-E Miscellaneous
Multi-robot navigation: This work focuses on single-robot navigation. Modeling robot-robot interactions in multi-robot navigation in crowded and constrained environments is an interesting future direction. To enable this, the extra robots and the interactions among all robots can be treated as extra nodes and spatial edges in the st-graph. Following our method, from the st-graph, robot-robot interactions can be modeled with self-attention or other methods, which serves as another submodule in each robot’s policy network.
Social awareness: The attention-based interaction framework presents a promising foundation for some social behaviors (e.x. treating humans and obstacles differently). However, HEIGHT focuses on optimizing task success and efficiency, and thus lacks explicit mechanisms and evaluation metrics for social awareness [francis2025principles]. Future work could integrate social norm as constraints or auxiliary objectives in RL training, as well as learning from human demonstrations or preference-based feedback. User studies can evaluate social awareness.